OpenStudy (anonymous):

a weightless rod is acted upon by parallel force of 4 N and 2 N at the ends .length of rod is 3 m . to keep rod in equilibrium 6 N should be applied where

rvc (rvc):



OpenStudy (michele_laino):

hint: the vector sum of the moments of the applied forces, has to be equal to zero: \[\Large {\mathbf{OO \times }}{{\mathbf{F}}_{\mathbf{1}}}{\mathbf{ + OP \times }}{{\mathbf{F}}_{\mathbf{2}}}{\mathbf{ + OQ \times }}{{\mathbf{F}}_{\mathbf{3}}}{\mathbf{ = 0}}\]  in this way the resultant of the applied forces is the null vector, so the first equation of rigid bodies is checked